# Content Bias in Deep Learning Image Age Approximation: A new Approach Towards better Explainability

### v2

@@ 2.1 Average Images
Refer to caption (a) class 1
Refer to caption (b) class 2
@@
Refer to caption $\overline{Y^{k}}$
Refer to caption $\overline{Y^{k}}_{c}$
Refer to caption $\overline{Y^{k}}_{r}$
Refer to caption $\overline{Y^{k}}_{f}$
@@ 2.2 Age Signal Embedded in Synthetic Images
Refer to caption (a) example $Y_{SYN}^{1}$
Refer to caption (b) example $Y_{SYN}^{2}$
Refer to caption (c) example mirroring
@@ 3.1 Evaluated Models
The first model considered is one that has already been used for image age classification, i.e., the AlexNet in transfer learning mode proposed in [6]. In [6], an image was cropped into 48 non-overlapping blocks of size $500\times 500$ and each block was fed separately through the network. The final class prediction was determined by a majority vote of each block’s predictions. It is not clearly described how the 48 non-overlapping blocks used are extracted, or how many epochs were trained with which batch size. In this work, the first 48 adjacent blocks, starting in the top left corner of the image, are extracted. When applying transfer learning over 60 epochs with a batch size of 4, the results reported in [6] can be confirmed. As defined in [6], the AdaMax is used as an optimizer with a learning rate of 0.0001.
@@
In [16], Bayar et al. proposed a preprocessing layer (the constrained convolutional layer) that should learn residuals. To enforce this layer to learn residuals, the following constraints
@@
are applied to each of the $K$ filters in the constraint convolution layer, where $k$ denotes the $k^{th}$ filter and the central weight $w$ is denoted by spatial index (0,0). Since an RGB input is considered in this work and the magnitude of the age signal may vary between color channels, the described constraint is applied to each layer of the filter kernel. The SRNet equipped with constrained convolutional layer (content suppression layer that increases the signal-to-noise ratio) is considered as fourth model and is denoted by SRNet-cs. All three SRNet variants (SRNet, SRNet-res and SRNet-cs) are trained in the ‘five-crop-fusion’ approach over a fixed number of 180 epochs with a batch size of 4 as done in [7, 8, 9]. The initial learning rate of 0.001 is decreased after 100 epochs to 0.0001 and the AdaMax is used as an optimizer.
The last model is again a model proposed in the context of image steganalysis, the ZhuNet [^1] https://github.com/1204BUPT/Zhu-Net-image-steganalysis [17]. The ZhuNet already includes a preprocessing layer. In particular, 30 basic high-pass filters of the Spatial Rich Model (SRM) [18] were used to extract residuals. This preprocessing layer can also be considered as content suppression layer that increases the signal-to-noise ratio. However, the parameters of the preprocessing layer (SRM filters) are not updated during training. For the ZhuNet, the defined optimizer is Stochastic Gradient Descend with a learning rate of 0.001 and momentum 0.95. The learning rate is divided by 5 after 20, 35, 50 and 65 epochs. A total of 80 epochs are conducted using the ‘five-crop-fusion’ approach. For synthetic images, the ZhuNet is trained over 240 epochs, as it takes longer to converge. For all five models, cross entropy is used as loss type.
@@ 3.2 Dataset
In the context of image age approximation, the NTIF database is used in [6, 8, 9, 22]. The authors in [6] divided the first 25 time-slots into five age classes. As already mentioned in the introduction, an overall classification accuracy of up to 88% is reported for these five classes (by AlexNet with transfer learning mode). To train the AlexNet, the same class definitions as in [6] are used (5 age classes). For training the SRNet variants and the ZhuNet only the first and last classes (i.e., time-slot 1-5 and 21-25, respectively) are considered (binary classification problem as done in [8, 9]). Overall, there is a time difference of about 4 months between time-slot 1-5 and 21-25.
For training the SRNet variants and the ZhuNet, images from two of our own imagers, a Canon PowerShotA720IS (Pc01) and a Nikon E7600 (Pn01), are used additionally. For both imagers, two classes are considered, with a time difference of about 12 and 15 years, respectively. Images from these two devices are also used in [8, 9, 3, 7]. The two sets of DFIs used to generate the synthetic images were also acquired with these two imagers. However, these two sets were not taken at the same time as the regular scene images.
@@ 3.4 Experimental Results
Tables 1, 2, 3, 4 and 5 show the results obtained when applying the proposed XAI method. As expected, if age classification is possible based on original input (i.e., $\overline{acc^{S}}>0.5$), age classification based on average images ($\overline{Y}$) and the structural components of the average images ($\overline{Y}_{r}$) is equally possible (i.e., $\overline{\Delta}(.)\approx 0$). However, if median filter residuals are used for training (SRNet-res, Table 3), $\overline{Y}_{r}$ seems to yield more stable results. Age classification based on average color ($\overline{Y}_{c}$) and the filtered average image ($\overline{Y}_{f}$) is not possible (i.e., $\overline{acc^{S}}-\overline{\Delta}(.)\approx 0.5$). This is reasonable since most likely in-field sensor defects are exploited by the network. Based on the results, it can be stated that the proposed XAI method works as expected.
@@
Table 1: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, when the AlexNet is trained on synthetic images (2 classes).
IDs $\overline{Y}$ $\overline{Y}_{c}$ $\overline{Y}_{r}$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
Table 2: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, when the SRNet is trained on synthetic images (2 classes).
IDs $\overline{Y}$ $\overline{Y}_{c}$ $\overline{Y}_{r}$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
Table 3: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, when the SRNet-res is trained on synthetic images (2 classes).
IDs $\overline{Y}$ $\overline{Y}_{c}$ $\overline{Y}_{r}$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
Table 4: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, when the SRNet-cs is trained on synthetic images (2 classes).
IDs $\overline{Y}$ $\overline{Y}_{c}$ $\overline{Y}_{r}$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
Table 5: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, when the ZhuNet is trained on synthetic images (2 classes).
IDs $\overline{Y}$ $\overline{Y}_{c}$ $\overline{Y}_{r}$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
Table 6: $\overline{\Delta}_{GEN}(\cdot)$, when the AlexNet is trained on regular scene images (5 classes).
IDs $\overline{Y}$ $\overline{Y}_{c}$ $\overline{Y}_{r}$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
Table 7: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, when the SRNet is trained on regular scene images (2 classes).
IDs $\overline{Y}$ $\overline{Y}_{c}$ $\overline{Y}_{r}$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
Nc01 0.26/0.39 0.39/0.39 0.27/0.58 0.38/0.48 0.96
Nc02 0.47/0.47 0.41/0.41 0.47/0.47 0.47/0.47 0.97
Nf01 0.10/0.10 0.28/0.28 0.35/0.35 0.19/0.19 0.85
Nf02 0.27/0.27 0.31/0.31 0.26/0.26 0.29/0.29 0.82
Nn01 0.30/0.30 0.39/0.39 0.39/0.39 0.39/0.39 0.89
Nn02 -0.01/-0.01 0.06/0.06 0.36/0.36 0.02/0.02 0.86
Np01 0.44/0.44 0.44/0.44 0.21/0.21 0.44/0.44 0.94
Np02 0.48/0.48 0.48/0.48 0.10/0.10 0.48/0.48 0.98
Ns01 0.25/0.25 0.28/0.28 0.38/0.38 0.25/0.25 0.88
Ns02 0.34/0.34 0.34/0.34 0.31/0.31 0.34/0.34 0.84
Pn01 0.16/0.16 0.16/0.16 0.33/0.33 0.16/0.16 0.83
Pc01 0.18/0.20 0.15/0.23 0.25/0.25 0.18/0.20 0.75
$\varnothing$ 0.27/0.28 0.31/0.31 0.31/0.33 0.30/0.31 0.88
Table 8: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, when the SRNet-res is trained on regular scene images (2 classes).
IDs $\overline{Y}$ $\overline{Y}_{c}$ $\overline{Y}_{r}$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
Nc01 0.33/0.46 0.23/0.35 0.20/0.33 0.19/0.31 0.92
Nc02 0.33/0.41 0.29/0.49 0.21/0.31 0.28/0.36 0.95
Nf01 0.40/0.40 0.40/0.40 0.21/0.38 0.34/0.47 0.91
Nf02 0.40/0.41 0.41/0.41 0.21/0.30 0.29/0.41 0.91
Nn01 0.43/0.43 0.43/0.43 0.32/0.32 0.33/0.35 0.93
Nn02 0.45/0.45 0.32/0.32 0.14/0.15 0.01/0.01 0.95
Np01 0.35/0.35 0.15/0.15 0.06/0.11 0.24/0.24 0.95
Np02 0.29/0.39 0.24/0.55 0.28/0.31 0.32/0.38 0.96
Ns01 0.43/0.43 0.39/0.48 0.13/0.13 0.35/0.35 0.93
Ns02 0.37/0.37 0.42/0.43 0.33/0.33 0.23/0.43 0.93
Pn01 0.37/0.37 0.24/0.34 0.32/0.43 0.27/0.36 0.91
Pc01 0.41/0.41 0.41/0.41 0.29/0.29 0.41/0.41 0.91
$\varnothing$ 0.38/0.41 0.33/0.40 0.23/0.28 0.27/0.34 0.93
Table 9: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, when the SRNet-cs is trained on regular scene images (2 classes).
IDs $\overline{Y}$ $\overline{Y}_{c}$ $\overline{Y}_{r}$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
Nc01 0.19/0.19 0.47/0.47 0.19/0.19 0.44/0.49 0.97
Nc02 0.26/0.26 0.47/0.47 0.27/0.27 0.47/0.47 0.97
Nf01 0.27/0.27 0.45/0.45 0.35/0.35 0.45/0.45 0.95
Nf02 0.11/0.11 0.45/0.45 0.20/0.20 0.40/0.40 0.95
Nn01 0.37/0.37 0.46/0.46 0.45/0.45 0.46/0.46 0.96
Nn02 0.37/0.37 0.46/0.46 0.43/0.43 0.46/0.46 0.96
Np01 0.06/0.06 0.46/0.46 0.07/0.07 0.43/0.47 0.96
Np02 0.20/0.20 0.47/0.47 0.11/0.11 0.37/0.41 0.97
Ns01 0.04/0.04 0.46/0.46 0.24/0.24 0.43/0.43 0.96
Ns02 0.38/0.38 0.46/0.46 0.41/0.41 0.46/0.46 0.96
Pn01 0.25/0.25 0.43/0.43 0.29/0.29 0.43/0.43 0.93
Pc01 0.14/0.14 0.43/0.43 0.13/0.13 0.38/0.38 0.93
$\varnothing$ 0.22/0.22 0.45/0.45 0.26/0.26 0.43/0.44 0.95
Table 10: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, when the ZhuNet is trained on regular scene images (2 classes).
IDs $\overline{Y}$ $\overline{Y}_{c}$ $\overline{Y}_{r}$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
Nc01 0.41/0.42 0.36/0.47 0.41/0.41 0.36/0.46 0.91
Nc02 0.42/0.42 0.42/0.42 0.42/0.42 0.42/0.42 0.92
Nf01 0.36/0.36 0.36/0.38 0.31/0.42 0.33/0.34 0.87
Nf02 0.24/0.29 0.28/0.28 0.32/0.32 0.23/0.23 0.82
Nn01 0.30/0.30 0.35/0.35 0.37/0.37 0.31/0.31 0.87
Nn02 0.14/0.14 0.30/0.36 0.38/0.38 0.16/0.16 0.94
Np01 0.45/0.45 0.45/0.45 0.45/0.45 0.45/0.45 0.95
Np02 0.44/0.44 0.43/0.43 0.39/0.49 0.44/0.44 0.94
Ns01 0.25/0.47 0.26/0.51 0.42/0.42 0.25/0.47 0.92
Ns02 0.36/0.36 0.40/0.40 0.40/0.40 0.39/0.39 0.90
Pn01 0.24/0.37 0.33/0.33 0.33/0.33 0.24/0.37 0.83
Pc01 0.29/0.29 0.29/0.41 0.36/0.36 0.32/0.32 0.86
$\varnothing$ 0.33/0.36 0.35/0.40 0.38/0.40 0.33/0.36 0.89
@@
The results for the AlexNet are reported in Table 6. With an average $\overline{\Delta}_{GEN}(\overline{Y})$ of 0.69 and an average $\overline{acc^{S}}$ of 0.94 (Table 6 last row), the average prediction accuracy of $\overline{Y}$ is close to the random prediction border (i.e., $0.94-0.69=0.25\approx 0.2$), indicating a strong influence of image content on the model inference. When the SRNet is trained on images of Nn02, it looks like an age signal is learned, as $\overline{\Delta}(\overline{Y})$ is even negative (Table 7 red values). However, the accuracy based on the average color ($\overline{Y}_{c}$) is similarly high. Since there is basically no age signal present in the average color, similar accuracy with the average color indicates that image content is exploited in this case as well. Across all evaluated imagers (Table 7 last row) the classification accuracy is similar for all average image variants.
@@
It is noticeable that age classification in the SRNet-res variant (Table 8 blue values) works better with average color ($\overline{Y}_{c}$) than with average images ($\overline{Y}$). This is surprising, since the color is strongly suppressed in median filter residuals. As observed with the baseline results, $\overline{Y}_{r}$ seems to yield more stable results with the SRNet-res variant (magenta value). Compared to the variants and imagers evaluated, the SRNet-cs variant when trained on images of Np01 and Ns01 (Table 9 green values) seems most likely to learn an age signal (i.e., a low value of $\overline{\Delta}(\overline{Y})$ combined with a high value of $\overline{\Delta}(\overline{Y}_{c})$). In both cases also high-frequency parts of the average image are relevant (i.e., high values of $\overline{\Delta}(\overline{Y}_{f})$). This could indicate that probably in-field sensor defects are exploited. Even when looking at the results across all devices (Table 9 last row) with an average $\overline{\Delta}(\overline{Y})$ of 0.22 and average $\overline{acc^{S}}$ of 0.95, the average prediction accuracy of $\overline{Y}$ is clearly above the prediction by chance border (i.e., $0.95-0.22=0.73>0.5$), while prediction based on $\overline{Y}_{c}$ or $\overline{Y}_{f}$ does not work. This would indicate that at least part of the inference is based on an age signal. In contrast, considering the ZhuNet (Table 10), the overall age classification accuracy of $\overline{Y}$ is close to the random prediction, indicating a strong influence of the image content.
@@
When comparing $\overline{\Delta}(\cdot)$ and $\overline{\Delta}_{GEN}(\cdot)$ a significant difference can only be observed with the model SRNet-res trained on regular scene images. In total, 11% of all evaluted runs (4 average image variants $\times$ 12 imagers $\times$ 8 runs) have an accuracy of less than 0.5. This is also significantly higher than for SRNet, SRNet-cs and ZhuNet with values of 3%, 2% and 5% respectively. When the models are trained on synthetic images (robust detection of an age signal), no accuracy values of less than 0.5 are observed.

### v3

@@ 2.1 Average Images
(a) class 1
(b) class 2
@@
$\overline{Y^{k}}$
$\overline{Y^{k}}_{c}$
$\overline{Y^{k}}_{r}$
$\overline{Y^{k}}_{f}$
@@ 2.2 Age Signal Embedded in Synthetic Images
(a) example $Y_{SYN}^{1}$
(b) example $Y_{SYN}^{2}$
(c) example mirroring
@@ 3.1 Evaluated Models
The first model considered is one that has already been used for image age classification, i.e., the AlexNet in transfer learning mode proposed in [6]. In [6], an image was cropped into 48 non-overlapping blocks of size $500\times 500$ and each block was fed separately through the network. The final class prediction was determined by a majority vote of each block’s predictions. It is not clearly described how the 48 non-overlapping blocks used are extracted, or how many epochs were trained with which batch size. In this work, the first 48 adjacent blocks, starting in the top left corner of the image, are extracted. When applying transfer learning over 60 epochs with a batch size of 4, the results reported in [6] can be confirmed. The AdaMax is used as an optimizer with a learning rate (lr) of 0.0001.
@@
In [16], Bayar et al. proposed a preprocessing layer (the constrained convolutional layer) that should learn residuals. To learn residuals, the following constraints
@@
are applied to each of the $K$ filters in the constraint convolution layer, where $k$ denotes the $k^{th}$ filter and the central weight $w$ is denoted by spatial index (0,0). Since an RGB input is considered in this work and the magnitude of the age signal may vary between color channels, the described constraint is applied to each layer of the filter kernel. The SRNet equipped with constrained convolutional layer (content suppression layer that increases the signal-to-noise ratio) is considered as fourth model and is denoted by SRNet-cs. All SRNet variants (SRNet, SRNet-res and SRNet-cs) are trained in the ‘five-crop-fusion’ approach over a fixed number of 180 epochs with a batch size of 4 as done in [7, 8, 9]. The lr of 0.001 is decreased to 0.0001 after 100 epochs and the AdaMax is used as an optimizer.
The last model is again a model proposed in the context of image steganalysis, the ZhuNet [^1] https://github.com/1204BUPT/Zhu-Net-image-steganalysis [17]. The ZhuNet already includes a preprocessing layer. In particular, 30 basic high-pass filters of the Spatial Rich Model (SRM) [18] were used to extract residuals. This preprocessing layer can also be considered as content suppression layer that increases the signal-to-noise ratio. The parameters of the preprocessing layer (SRM filters) are not updated during training. For the ZhuNet, the defined optimizer is Stochastic Gradient Descend with a lr of 0.001 and momentum 0.95. The lr is divided by 5 after 20, 35, 50 and 65 epochs. A total of 80 epochs are conducted using the ‘five-crop-fusion’ approach. For synthetic images, the ZhuNet is trained over 240 epochs, as it takes longer to converge. For all five models, cross entropy is used as loss type. A comparison of SRNet and ZhuNet is given in [17].
@@ 3.2 Dataset
In the context of image age approximation, the NTIF database is used in [6, 8, 9, 22]. The authors in [6] divided the first 25 time-slots into five age classes. As already mentioned in the introduction, an overall classification accuracy of up to 88% is reported for these five classes (by AlexNet with transfer learning mode). To train the AlexNet, the same class definitions as in [6] are used (5 age classes). For training the SRNet variants and the ZhuNet only the first and last classes (i.e., time-slot 1-5 and 21-25, respectively) are considered (binary classification problem as done in [8, 9]). There is a time difference of about 4 months between time-slot 1-5 and 21-25. For training the SRNet variants and the ZhuNet, images from two of our own imagers, a Canon PowerShotA720IS (Pc01) and a Nikon E7600 (Pn01), are used additionally. For both imagers, two classes are considered, with a time difference of about 12 and 15 years, respectively. Images from these two devices are also used in [8, 9, 3, 7]. The two sets of DFIs used to generate the synthetic images were also acquired with these two imagers. However, these two sets were not taken at the same time as the regular scene images.
Between the first and second class of Pc01, a total of 27 new ‘strong’ in-field sensor defects have developed, 3 of which are located in the five-crop regions. These ‘strong’ defects are detected by thresholding the median filter residuals of the generated average images. For all other imagers, no ‘strong’ in-field sensor defects are detected in the five-crop regions. It is also noticeable that the camera parameters (ISO, exposure time and focal length) differ significantly between the classes (e.g., the mean ISO/exp.time/focal length for class 1 and 2 of Nc02 are 342/0.013/7.45 and 163/0.0076/5 with a standard deviation of 360/0.020/4.84 and 89/0.011/0), which changes the image characteristics (i.e., a higher ISO/exp.time lead to more noise, while different focal lengths can lead to different distortions). This is also an indication of different lighting conditions between the two classes.
@@ 3.4 Experimental Results
Table 1 shows the results obtained when applying the proposed XAI method. As expected, if age classification is possible based on original input (i.e., $\overline{acc^{S}}>0.5$), age classification based on average images ($\overline{Y}$) and the structural components of the average images ($\overline{Y}_{r}$) is equally possible (i.e., $\overline{\Delta}(.)\approx 0$). However, if median filter residuals are used for training (SRNet-res, Table 1), $\overline{Y}_{r}$ seems to yield more stable results. Age classification based on average color ($\overline{Y}_{c}$) and the filtered average image ($\overline{Y}_{f}$) is not possible (i.e., $\overline{acc^{S}}-\overline{\Delta}(.)\approx 0.5$). This is reasonable since most likely in-field sensor defects are exploited by the network. Based on the results, it can be stated that the proposed XAI method works as expected.
@@
Table 1: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, models trained on syn. imgs (2 classes). Recall: An age signal is present in $\overline{Y}$ and $\overline{Y}_{r}$ and not in $\overline{Y}_{c}$. Expectation if an age signal is learned, $\downarrow:\overline{\Delta}\approx 0$ and $\uparrow:\overline{\Delta}>>0$. $\overline{Y}_{f}$, high frequency image components (e.g., defects) are filtered out.
AlexNet
IDs $\overline{Y}\downarrow$ $\overline{Y}_{c}\uparrow$ $\overline{Y}_{r}\downarrow$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
SRNet
IDs $\overline{Y}\downarrow$ $\overline{Y}_{c}\uparrow$ $\overline{Y}_{r}\downarrow$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
SRNet-res
IDs $\overline{Y}\downarrow$ $\overline{Y}_{c}\uparrow$ $\overline{Y}_{r}\downarrow$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
SRNet-cs
IDs $\overline{Y}\downarrow$ $\overline{Y}_{c}\uparrow$ $\overline{Y}_{r}\downarrow$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
ZhuNet
IDs $\overline{Y}\downarrow$ $\overline{Y}_{c}\uparrow$ $\overline{Y}_{r}\downarrow$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
Table 2: $\overline{\Delta}_{GEN}(\cdot)$, AlexNet trained on reg. scene imgs (5 classes). Recall: An age signal is present in $\overline{Y}$ and $\overline{Y}_{r}$ and not in $\overline{Y}_{c}$. Expectation if an age signal is learned, $\downarrow:\overline{\Delta}\approx 0$ and $\uparrow:\overline{\Delta}>>0$. $\overline{Y}_{f}$, high frequency image components (e.g., defects) are filtered out.
IDs $\overline{Y}\downarrow$ $\overline{Y}_{c}\uparrow$ $\overline{Y}_{r}\downarrow$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
@@
Table 3: $\overline{\Delta}(\cdot)/\overline{\Delta}_{GEN}(\cdot)$, models trained on reg. scene imgs (2 classes). Recall: An age signal is present in $\overline{Y}$ and $\overline{Y}_{r}$ and not in $\overline{Y}_{c}$. Expectation if an age signal is learned, $\downarrow:\overline{\Delta}\approx 0$ and $\uparrow:\overline{\Delta}>>0$. $\overline{Y}_{f}$, high frequency image components (e.g., defects) are filtered out.
SRNet SRNet-res
IDs $\overline{Y}\downarrow$ $\overline{Y}_{c}\uparrow$ $\overline{Y}_{r}\downarrow$ $\overline{Y}_{f}$ $\overline{acc^{S}}$ IDs $\overline{Y}\downarrow$ $\overline{Y}_{c}\uparrow$ $\overline{Y}_{r}\downarrow$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
Nc01 0.26/0.39 0.39/0.39 0.27/0.58 0.38/0.48 0.96 Nc01 0.33/0.46 0.23/0.35 0.20/0.33 0.19/0.31 0.92
Nc02 0.47/0.47 0.41/0.41 0.47/0.47 0.47/0.47 0.97 Nc02 0.33/0.41 0.29/0.49 0.21/0.31 0.28/0.36 0.95
Nf01 0.10/0.10 0.28/0.28 0.35/0.35 0.19/0.19 0.85 Nf01 0.40/0.40 0.40/0.40 0.21/0.38 0.34/0.47 0.91
Nf02 0.27/0.27 0.31/0.31 0.26/0.26 0.29/0.29 0.82 Nf02 0.40/0.41 0.41/0.41 0.21/0.30 0.29/0.41 0.91
Nn01 0.30/0.30 0.39/0.39 0.39/0.39 0.39/0.39 0.89 Nn01 0.43/0.43 0.43/0.43 0.32/0.32 0.33/0.35 0.93
Nn02 -0.01/-0.01 0.06/0.06 0.36/0.36 0.02/0.02 0.86 Nn02 0.45/0.45 0.32/0.32 0.14/0.15 0.01/0.01 0.95
Np01 0.44/0.44 0.44/0.44 0.21/0.21 0.44/0.44 0.94 Np01 0.35/0.35 0.15/0.15 0.06/0.11 0.24/0.24 0.95
Np02 0.48/0.48 0.48/0.48 0.10/0.10 0.48/0.48 0.98 Np02 0.29/0.39 0.24/0.55 0.28/0.31 0.32/0.38 0.96
Ns01 0.25/0.25 0.28/0.28 0.38/0.38 0.25/0.25 0.88 Ns01 0.43/0.43 0.39/0.48 0.13/0.13 0.35/0.35 0.93
Ns02 0.34/0.34 0.34/0.34 0.31/0.31 0.34/0.34 0.84 Ns02 0.37/0.37 0.42/0.43 0.33/0.33 0.23/0.43 0.93
Pn01 0.16/0.16 0.16/0.16 0.33/0.33 0.16/0.16 0.83 Pn01 0.37/0.37 0.24/0.34 0.32/0.43 0.27/0.36 0.91
Pc01 0.18/0.20 0.15/0.23 0.25/0.25 0.18/0.20 0.75 Pc01 0.41/0.41 0.41/0.41 0.29/0.29 0.41/0.41 0.91
$\varnothing$ 0.27/0.28 0.31/0.31 0.31/0.33 0.30/0.31 0.88 $\varnothing$ 0.38/0.41 0.33/0.40 0.23/0.28 0.27/0.34 0.93
SRNet-cs ZhuNet
IDs $\overline{Y}\downarrow$ $\overline{Y}_{c}\uparrow$ $\overline{Y}_{r}\downarrow$ $\overline{Y}_{f}$ $\overline{acc^{S}}$ IDs $\overline{Y}\downarrow$ $\overline{Y}_{c}\uparrow$ $\overline{Y}_{r}\downarrow$ $\overline{Y}_{f}$ $\overline{acc^{S}}$
Nc01 0.19/0.19 0.47/0.47 0.19/0.19 0.44/0.49 0.97 Nc01 0.41/0.42 0.36/0.47 0.41/0.41 0.36/0.46 0.91
Nc02 0.26/0.26 0.47/0.47 0.27/0.27 0.47/0.47 0.97 Nc02 0.42/0.42 0.42/0.42 0.42/0.42 0.42/0.42 0.92
Nf01 0.27/0.27 0.45/0.45 0.35/0.35 0.45/0.45 0.95 Nf01 0.36/0.36 0.36/0.38 0.31/0.42 0.33/0.34 0.87
Nf02 0.11/0.11 0.45/0.45 0.20/0.20 0.40/0.40 0.95 Nf02 0.24/0.29 0.28/0.28 0.32/0.32 0.23/0.23 0.82
Nn01 0.37/0.37 0.46/0.46 0.45/0.45 0.46/0.46 0.96 Nn01 0.30/0.30 0.35/0.35 0.37/0.37 0.31/0.31 0.87
Nn02 0.37/0.37 0.46/0.46 0.43/0.43 0.46/0.46 0.96 Nn02 0.14/0.14 0.30/0.36 0.38/0.38 0.16/0.16 0.94
Np01 0.06/0.06 0.46/0.46 0.07/0.07 0.43/0.47 0.96 Np01 0.45/0.45 0.45/0.45 0.45/0.45 0.45/0.45 0.95
Np02 0.20/0.20 0.47/0.47 0.11/0.11 0.37/0.41 0.97 Np02 0.44/0.44 0.43/0.43 0.39/0.49 0.44/0.44 0.94
Ns01 0.04/0.04 0.46/0.46 0.24/0.24 0.43/0.43 0.96 Ns01 0.25/0.47 0.26/0.51 0.42/0.42 0.25/0.47 0.92
Ns02 0.38/0.38 0.46/0.46 0.41/0.41 0.46/0.46 0.96 Ns02 0.36/0.36 0.40/0.40 0.40/0.40 0.39/0.39 0.90
Pn01 0.25/0.25 0.43/0.43 0.29/0.29 0.43/0.43 0.93 Pn01 0.24/0.37 0.33/0.33 0.33/0.33 0.24/0.37 0.83
Pc01 0.14/0.14 0.43/0.43 0.13/0.13 0.38/0.38 0.93 Pc01 0.29/0.29 0.29/0.41 0.36/0.36 0.32/0.32 0.86
$\varnothing$ 0.22/0.22 0.45/0.45 0.26/0.26 0.43/0.44 0.95 $\varnothing$ 0.33/0.36 0.35/0.40 0.38/0.40 0.33/0.36 0.89
@@
The results for the AlexNet are reported in Table 2. With an average $\overline{\Delta}_{GEN}(\overline{Y})$ of 0.69 and an average $\overline{acc^{S}}$ of 0.94 (Table 2 last row), the average prediction accuracy of $\overline{Y}$ is close to the random prediction border (i.e., $0.94-0.69=0.25\approx 0.2$), indicating a strong influence of image content on the model inference. When the SRNet is trained on images of Nn02, it looks like an age signal is learned, as $\overline{\Delta}(\overline{Y})$ is even negative (Table 3, SRNet, bold values). However, the accuracy based on the average color ($\overline{Y}_{c}$) is similarly high. Since there is basically no age signal present in the average color, similar accuracy with the average color indicates that image content is exploited in this case as well. Across all evaluated imagers (Table 3, SRNet, last row) the classification accuracy is similar for all average image variants.
@@
It is noticeable that age classification in the SRNet-res variant (Table 3, SRNet-res, bold values) works better with average color ($\overline{Y}_{c}$) than with average images ($\overline{Y}$). This is surprising, since the color is strongly suppressed in median filter residuals. As observed with the baseline results, $\overline{Y}_{r}$ seems to yield more stable results with the SRNet-res variant. Compared to the variants and imagers evaluated, the SRNet-cs variant when trained on images of Np01 and Ns01 (Table 3, SRNet-cs, bold values) seems most likely to learn an age signal (i.e., a low value of $\overline{\Delta}(\overline{Y})$ combined with a high value of $\overline{\Delta}(\overline{Y}_{c})$). In both cases also high-frequency parts of the average image are relevant (i.e., high values of $\overline{\Delta}(\overline{Y}_{f})$). This could indicate that probably in-field sensor defects are exploited. Even when looking at the results across all devices (Table 3, SRNet-cs, last row) with an average $\overline{\Delta}(\overline{Y})$ of 0.22 and average $\overline{acc^{S}}$ of 0.95, the average prediction accuracy of $\overline{Y}$ is clearly above the prediction by chance border (i.e., $0.95-0.22=0.73>0.5$), while prediction based on $\overline{Y}_{c}$ or $\overline{Y}_{f}$ does not work. This would indicate that at least part of the inference is based on an age signal. In contrast, considering the ZhuNet (Table 3, ZhuNet), the overall age classification accuracy of $\overline{Y}$ is close to the random prediction, indicating a strong influence of the image content.
@@
When comparing $\overline{\Delta}(\cdot)$ and $\overline{\Delta}_{GEN}(\cdot)$ a significant difference can only be observed with the model SRNet-res trained on regular scene images. In total, 11% of all evaluated runs (4 average image variants $\times$ 12 imagers $\times$ 8 runs) have an accuracy of less than 0.5. This is also significantly higher than for SRNet, SRNet-cs and ZhuNet with values of 3%, 2% and 5% respectively. When the models are trained on synthetic images (robust detection of an age signal), no accuracy values of less than 0.5 are observed.